# InfoLossQA: Characterizing and Recovering Information Loss in Text Simplification

### v1

@@ Abstract
Text simplification aims to make technical texts more accessible to laypeople but often results in deletion of information and vagueness. This work proposes InfoLossQA, a framework to characterize and recover simplification-induced information loss in form of question-and-answer (QA) pairs. Building on the theory of Question Under Discussion, the QA pairs are designed to help readers deepen their knowledge of a text. We conduct a range of experiments with this framework. First, we collect a dataset of 1,000 linguist-curated QA pairs derived from 104 LLM simplifications of scientific abstracts of medical studies. Our analyses of this data reveal that information loss occurs frequently, and that the QA pairs give a high-level overview of what information was lost. Second, we devise two methods for this task: end-to-end prompting of open-source and commercial language models, and a natural language inference pipeline. With a novel evaluation framework considering the correctness of QA pairs and their linguistic suitability, our expert evaluation reveals that models struggle to reliably identify information loss and applying similar standards as humans at what constitutes information loss.
@@
[…]
@@ 1 Introduction
Technical texts, many of which exist in high-stake domains (e.g., medicine), are often written in a language incomprehensible to lay readers. Improving the accessibility of texts may help address wider social issues such as disinformation UNSDG (2021) and improve access to higher education Goff et al. (2004). Automatic text simplification that rewrites text into plain language may therefore be a technology for good. With the adoption of LLMs, document-level text simplification has significantly progressed in recent years August et al. (2023); Laban et al. (2023); Agrawal and Carpuat (2023).
Simplification is an inherently lossy process: Even when done by professional editors, the resulting plain language tends to lose details and become more generic Li and Nenkova (2015), and some content is omitted Zhong et al. (2020). The over-simplification of content—including excessive deletion and vagueness—may lead to misinterpretations and factual errors Devaraj et al. (2022), or reduced comprehension Agrawal and Carpuat (2023). Therefore, we consider the following question: How can we characterize information loss and help readers recover what is lost in an intuitive and understandable manner?
@@
This paper presents a question generation and answering framework to characterize and recover information loss during simplification in a reader-centric manner. Inspired by the theoretical work of “potential questions” Onea (2016) and expectation-driven Questions Under Discussion (QUD, Kehler and Rohde, 2017), information loss is characterized in the form of pragmatically felicitous questions rooted in simplified texts, without assuming access to the original text. Information loss is then recovered by answering these questions in plain language, based on the original (technical) text (Figure 1). Informed by QUD, the questions are curiosity-driven and information-seeking, and are subject to well-defined linguistic constraints Riester et al. (2018); Wu et al. (2023). While prior work used factoid QAs to describe specific factual changes Cole et al. (2023), the QUD framework suggests higher-level questions, and allows us to characterize general yet subtle language changes pervasive in factually equivalent texts. These include lower specificity Louis and Nenkova (2011) and greater degrees of vagueness van Deemter (2012). We envision the resulting QA pairs being directly incorporated in interactive simplification tools Fok et al. (2023).
We first introduce InfoLossQA, a linguist-curated dataset grounded in the scenario of lay readers digesting simplified abstracts of randomized controlled trials (RCTs) in medicine and healthcare. We focus on RCTs because they combine several important and illustrative properties: They use highly technical/inaccessible language; lay audiences often having keen interest in their content; and misinterpretations are risky, given the implications for health. InfoLossQA consists of 1,000 QA pairs across 104 RCT abstracts simplified by GPT-4 (prior work has shown that GPT-3.5 is a capable simplifier of medical texts; Shaib et al., 2023). The QA pairs provide a rich description of what information is lost both on a pragmatic and semantic level. Our extensive analysis shows that technical sections of RCTs are particularly prone for information loss. Questions often elaborate about why and how a clinical trial was conducted, going beyond frequently asked questions previously identified in this domain August et al. (2023).
We then present two methods to perform the question generation and answering task. First, an end-to-end model, where we prompt open-source and API-based LLMs with detailed task instructions and a one-shot example. Second, a natural language inference (NLI) pipeline which reasons over entailment of atomic facts Stacey et al. (2023).
@@
To summarize, we make the following technical contributions:
  1. 1.
    We propose InfoLossQA, a novel task and linguist-curated dataset to characterize and reveal simplification-induced information loss in form of QA pairs.
  2. 2.
    We establish two baselines at this task, including end-to-end LLMs and an NLI pipeline.
  3. 3.
    We propose a comprehensive framework to evaluate models at this task and collect human quality judgments.
We released the experiment code and the dataset at https://InfoLossQA.ikim.nrw/.
@@ 2 The InfoLossQA Task
Information loss is inherent to simplification, insofar as the task demands producing outputs which are both comparatively uncomplicated and reasonably concise. Characterizing and revealing the specific content elided is critical to (1) provide a path for users to achieve better understanding than only seeing the simplified text, supporting users’ growth Gooding (2022), and (2) allow researchers to evaluate the quality of a simplified text, given that omitting or oversimplifying key information may yield misleading texts Devaraj et al. (2022). Moreover, we envision simplification as an interactive process in which a user may ask for additional details on particular points. InfoLossQA defines a task and dataset that entails generating QA pairs revealing to lay readers the information missing from a simplified text, compared to its original.
@@
Given a pair of documents, the original text $X_{\text{src}}$ and the simplified text $X_{\text{tgt}}$, the task entails generating QA pairs that each reveal to lay readers one instance of information loss. We define an instance of information loss as a tuple $(X_{\text{src}},X_{\text{tgt}},Q,A,E,R)$. A potential question $Q$ asks for information absent from $X_{\text{tgt}}$. The answer $A$ to the question $Q$ provides the missing information. We additionally require two forms of localization or grounding to support richer analysis and scrutiny of errors in the generated QA pairs. First, each answer $A$ must be supported by evidence $E\in X_{\text{src}}$ extracted from the original text. Second, each $Q$ and $A$ pair may be supported by an extractive rationale $R\in X_{\text{tgt}}$ which localizes where the information loss or vagueness occurs within the simplified text. Note that while $E$ always exists for an instance, $R$ may not (some information in $X_{\text{src}}$ may be completely absent from $X_{\text{tgt}}$). This formulation permits multiple information loss instances for a given $X_{\text{src}}$ and $X_{\text{tgt}}$ pair. We do not constrain the length or linguistic unit(s) of the extractive information loss spans ($E$ and $R$): they can comprise words, phrases, sentences, or entire paragraphs, and a single $E$ or $R$ can be one or more spans.
@@
    Oversimplification. Pieces of information that were simplified to the extent that they are vague or devoid of their original meaning. This is the case where the extractive rationale $R$ is identified.
These categories are meant to be flexible to capture a variety of information transformations, but also sufficiently concrete to be operationalized in an annotation protocol (Section 3).
@@
One of our primary goals is to generate QA pairs that might help users gain a fuller understanding of the original text Fok et al. (2023). The desire for a deeper understanding may arise when a user has a keen interest in aspects of the original text that go beyond what is presented in the simplified version. Enabling such interactions may also allow individuals to calibrate their trust in AI-simplified outputs McDermott and Brink (2019). We pose two requirements for the QA pairs:
@@
Example.
We illustrate the different elements and linguistic constraints of the task with an example. Consider the following passages taken from the abstract of a clinical study:
Original ($X_{\text{src}}$): These results indicate that acute/chronic endurance ($E_{1}$) arm-cranking with EMS applied to the lower limbs improves the brachial artery endothelial function ($E_{2}$) more markedly than the same exercise without EMS.
Simplified ($X_{\text{tgt}}$): The study concluded that doing the arm-cranking exercise with EMS on the lower body can improve arm function ($R_{2}$) more than doing the exercise without EMS.
This example demonstrates the challenging nature of the task. To identify information loss, models need to bridge two language levels to decide on the presence and nature of information loss. Relying on lexical overlap, ordering, or other surface level properties is infeasible. The example includes both a deletion (EMS helps in “acute/chronic” settings) and an oversimplification (“improve arm function” is too broad given that EMS improves “artery function”). By contrast, the lexical simplification “lower limbs” $\rightarrow$ “lower body” may be viewed as functionally equivalent, and an arguably unproblematic case of information loss.
This simplification gives rise to two QA pairs:
  1. Q1.
    On what timeframes does EMS improve training? A. It helps in single application (acute) or repeated application (chronic).
  2. Q2.
    How did the researchers measure how well EMS works? A. The researchers measure to what extent the main artery of the arm widens, which is called flow-mediated vasodilation (FMD).
Note that the QA pairs fulfill the stylistic and linguistic constraints we defined above: the technical material is explained as necessary, and the questions are interpretable for readers without having seen the original text fulfilling the givenness constraint. For contrast, an invalid (↯) question conflicting with the givenness constraint is as follows:
  1. Q3.
    Did EMS improve training in acute and chronic applications?
While Q1 and Q3 ask about the same information, Q1 would be more natural to lay readers because it could conceivably be asked without having seen the original text (it does not presuppose that the study looked at short-term and long-term effects).
@@ 3.1 Simplification Corpus
In this work, we focus on medical texts, which provide an important and illustrative case where lay individuals may be keenly interested in understanding newly published science. More specifically, we collect annotations on abstracts of medical journal articles that describe the conduct and findings of Randomized Controlled Trials (RCTs). RCTs are the foundation of evidence-based medicine Sackett (1998) and provide the critical comparative effectiveness findings that inform patient care. But articles describing RCTs use technical language, and so are effectively inaccessible to (lay) patients Zuccala (2010); August et al. (2023). This means that individuals may already interact with LLMs to simplify such texts. But automated simplification of RCTs may result in information loss, which in turn risks consumers misinterpreting findings.
To better quantify information loss in LLM-simplified outputs, we create a parallel simplification corpus with GPT-4 (gpt-4-0613). [^2] There is no existing curated parallel corpus for RCT simplification. Initially, we considered obtaining medical abstracts that are not RCTs with their associated Plain Language Summaries (PLS) from journals, but found that these PLS vary too significantly in content and style to form a parallel corpus, a problem also noted in prior related work Devaraj et al. (2021), partly due to the ongoing debate what an effective PLS should look like Jelicic Kadic et al. (2016). On the related task of plain language summarization of full RCT publications, Shaib et al. (2023) observed very low error rates (less than 10% of summaries affected), but we do not yet have a quantitative understanding of information loss in LLM-simplified RCT texts. We sample RCT abstracts from the Evidence Inference dataset DeYoung et al. (2020b); Lehman et al. (2019), originally sourced from PubMed. We provide complete generation parameters and example outputs in Appendix A.
@@ 3.2 Annotation Process
The process for annotating InfoLossQA is as follows. First, annotators examine both the original and simplified versions of the text. Next they identify and highlight any spans where information is missing from the simplified text. For each highlighted instance, annotators compose a QA-pair.
(a) Distribution of QA pairs over RCT abstract sections. Some QAs are localized in multiple sections (multi.) or in a rare section (other). We also report the percentage of QAs that falls into abstracts without section structure (19.2% of abstracts; no secs.).
(b) Number of QA pairs per sentence for a random selection of 25 documents. The x-axis are sentence indices.
Figure 2: Distribution of information loss over sections and sentences. Humans produce a similar distribution of questions by section (a), but the questions differ in their localization (b). Comparing humans to models, we can see significant differences where questions are asked, and by extension also what they are about.
@@
Technical texts like RCTs have a long tail of information that can be lost during simplification. For example, numerical results like confidence intervals and significance scores. Annotating all information loss is intractable, and the resulting QAs would likely overwhelm readers. Therefore, we grounded annotation in a scenario: annotators were asked to imagine a patient who could benefit from better understanding the RCT at hand, and instructed to focus on information that may be salient in this scenario.
@@
We hired three expert annotators majoring in linguistics with experience annotating medical texts. To ensure high annotation quality, we conducted a pilot phase with written feedback and held frequent review meetings to discuss questions. Annotators were paid \$15/hr. [^3] Higher than local minimum wage at the time of study. The median duration per document was 22 minutes. We used the Thresh toolkit Heineman et al. (2023) for annotation which allows parallel text highlights. See Appendix B for annotation guidelines and screenshots of the interface that we used.
@@
We collected annotations on 104 abstract-simplification pairs (average number of tokens: $|X_{\text{src}}|=312$; $|X_{\text{tgt}}|=271$) yielding 1,000 questions and answers (average number of tokens: $|Q|=11.4$; $|A|=30.3$). [^4] Tokenizing with nltk Bird et al. (2009). Each annotator wrote on average between 4.1 and 5.7 QA pairs (i.e., information loss instances) per document. We observe that oversimplification is more prevalent (74.1% of questions) in this corpus than complete deletion of information (25.9% of questions). A detailed breakdown of the statistics by annotator is given in Table A.2 and Section 3.3 analyzes to what extent annotators ask similar questions.
Question Type % All % O % D Examples
Procedural. Asking for the procedures, tools, or methods by which a certain outcome is achieved.
34.3 34.1 34.7
• How did they measure the patients’ opioid medication needs?
• How did the study control for bias?
Concept. Asking for a definition of an event or a concept.
25.7 28.1 18.9
• What kind of hip surgery were patients undergoing? • What type of mental illnesses are being studied?
Extent. Asking for the extent or quantity of an event or a concept.
17.0 19.2 10.8
• How many patients were in each group? • How much lurasidone was given to the patients in the trial?
Comparison. Asking for comparison among multiple events or concepts.
8.3 8.9 6.6
• How much did the abnormal blood vessels reduce in group 1 compared to group 2? • How did pre-meal insulin affect platelet activation compared to the placebo?
Consequence. Asking for the consequences or results of an event.
5.7 3.0 13.5
• What else does increased blood flow do? • What was the correlation between measures for total daily calcium intake?
Cause. Asking for the cause or reason for an event or a concept.
4.3 1.6 12.0
• What was a motivation behind the study? • Why are soy products being looked at in this study?
Example. Asking for example(s) or instance(s) of an event or a concept.
3.4 4.3 0.8
• What are the other brain functions that were tested? • What medication is typically used for outpatients?
Disjunction. Asking for the true one given multiple events or concepts, where comparison among options is not needed.
0.6 0.5 0.8
• Which one of the three components did this group’s supplement contain? • Which gel saw the increase in beneficial microbes after 24 hours?
Verification. Asking for the truthfulness of an event or a concept.
0.7 0.3 1.9
• Were there any other changes in blood sugar level observed in the study? • Were both eyes of each patient tested?
Judgmental. Asking for the opinions of the answerer’s own.
Not observed.

### v2

@@ Abstract
Text simplification aims to make technical texts more accessible to laypeople but often results in deletion of information and vagueness. This work proposes InfoLossQA, a framework to characterize and recover simplification-induced information loss in form of question-and-answer (QA) pairs. Building on the theory of Questions Under Discussion, the QA pairs are designed to help readers deepen their knowledge of a text. First, we collect a dataset of 1,000 linguist-curated QA pairs derived from 104 LLM simplifications of English medical study abstracts. Our analyses of this data reveal that information loss occurs frequently, and that the QA pairs give a high-level overview of what information was lost. Second, we devise two methods for this task: end-to-end prompting of open-source and commercial language models, and a natural language inference pipeline. With a novel evaluation framework considering the correctness of QA pairs and their linguistic suitability, our expert evaluation reveals that models struggle to reliably identify information loss and applying similar standards as humans at what constitutes information loss. [^1] Code, dataset and an interactive data viewer is available at https://jantrienes.github.io/ts-info-loss/.
@@
Jan Trienes1,2,4thanks: Work done while visiting UT Austin. Sebastian Joseph3 Jörg Schlötterer4,5 Christin Seifert4Kyle Lo6Wei Xu7Byron C. Wallace8Junyi Jessy Li3 1University of Duisburg-Essen 2Institute for AI in Medicine, University Hospital Essen 3The University of Texas at Austin 4University of Marburg 5University of Mannheim 6Allen Institute for AI 7Georgia Institute of Technology 8Northeastern University jan.trienes@uni-marburg.dejessy@utexas.edu
@@ 1 Introduction
Technical texts, many of which exist in high-stake domains (e.g., medicine), are often written in a language incomprehensible to lay readers. Improving the accessibility of such texts may help address wider social issues, e.g., disinformation UNSDG (2021) and access to higher education Goff et al. (2004). Automatic text simplification that rewrites text into plain language may therefore be a technology for good. With the adoption of LLMs, document-level text simplification has significantly progressed in recent years August et al. (2023); Laban et al. (2023); Agrawal and Carpuat (2024).
Simplification is an inherently lossy process: Even when done by professional editors, the resulting plain language tends to lose details and become more generic Li and Nenkova (2015), and some content is omitted Zhong et al. (2020). The over-simplification of content—including excessive deletion and vagueness—may lead to reduced comprehension Agrawal and Carpuat (2024) or in its worst case to misinterpretations and factual errors Devaraj et al. (2022). Therefore, we consider the following question: How can we characterize information loss and help readers recover what is lost in an intuitive and understandable manner?
@@
This paper presents a question generation and answering framework to characterize and recover information loss during simplification in a reader-centric way. Inspired by the theoretical work of “potential questions” Onea (2016) and expectation-driven Questions Under Discussion (QUD, Kehler and Rohde, 2017), information loss is characterized in the form of pragmatically felicitous questions rooted in simplified texts, without assuming access to the original text. Information loss is then recovered by answering these questions in plain language, based on the original (technical) text (Figure 1). Informed by QUD, the questions are curiosity-driven and information-seeking, and are subject to well-defined linguistic constraints Riester et al. (2018); Wu et al. (2023). While prior work used factoid QAs to describe specific factual changes Cole et al. (2023), the QUD framework suggests higher-level questions, and allows us to characterize general yet subtle language changes pervasive in factually equivalent texts. These include lower specificity Louis and Nenkova (2011) and greater degrees of vagueness van Deemter (2012). We envision the resulting QA pairs being directly incorporated in interactive simplification tools Fok et al. (2023).
Contribution 1: Task formulation and dataset.
We first introduce InfoLossQA, a linguist-curated dataset grounded in the scenario of lay readers digesting simplified abstracts of randomized controlled trials (RCTs) in medicine. We focus on RCTs because they combine important and illustrative properties: They use highly technical/inaccessible language; lay audiences often having keen interest in their content; and misinterpretations are risky in the healthcare domain. InfoLossQA consists of 1,000 QA pairs across 104 RCT abstracts simplified by GPT-4 (prior work has shown that GPT-3.5 is a capable simplifier of medical texts; Shaib et al., 2023). The QA pairs provide a rich description of what information is lost both on a pragmatic and semantic level. Our analyses show that questions often elaborate about why and how an RCT was conducted, going beyond frequently asked questions in this domain August et al. (2023).
Contribution 2: Automatic methods.
We then present two methods to perform the question generation and answering task. First, an end-to-end model, where we prompt several open-source and API-based LLMs with detailed task instructions and a one-shot example. Second, a natural language inference (NLI) pipeline which reasons over entailment of atomic facts Stacey et al. (2023).
Contribution 3: Evaluation framework and human quality judgments.
@@ 2 The InfoLossQA Task
Motivation.
Information loss is inherent to simplification, insofar as the task typically demands producing outputs which are both comparatively uncomplicated and reasonably concise. Characterizing and revealing the elided content is critical to (1) provide a path for users to achieve better understanding than only seeing the simplified text, supporting users’ growth Gooding (2022), and (2) to allow researchers to evaluate the quality of a simplified text, given that omitting or oversimplifying key information may yield misleading texts Devaraj et al. (2022).
@@
Given a pair of documents, the original text $X_{\text{src}}$ and the simplified text $X_{\text{tgt}}$, the task entails generating QA pairs that each reveal to lay readers one instance of information loss. We define an instance of information loss as a tuple $(X_{\text{src}},X_{\text{tgt}},Q,A,E,R)$. A potential question $Q$ asks for information absent from $X_{\text{tgt}}$. The answer $A$ provides the missing information. We additionally require two forms of localization or grounding to support richer analysis and scrutiny of errors in the generated QA pairs. First, each answer $A$ must be supported by evidence $E\in X_{\text{src}}$ extracted from the original text. Second, each QA pair may be supported by an extractive rationale $R\in X_{\text{tgt}}$ which localizes where the information loss or vagueness occurs within the simplified text. Note that while $E$ always exists, $R$ may not (some information in $X_{\text{src}}$ may be completely absent from $X_{\text{tgt}}$).
This formulation permits multiple information loss instances for a given $X_{\text{src}}$ and $X_{\text{tgt}}$ pair. We do not constrain the length or linguistic unit(s) of the extractive spans ($E$ and $R$): they can comprise words, phrases, sentences, or entire paragraphs, and a single $E$ or $R$ can be one or more spans.
@@
    Oversimplification. Pieces of information that were simplified to the extent that they are vague or devoid of their original meaning. This is where the rationale $R$ is identified.
These categories are meant to be flexible to capture a variety of information transformations, but also concrete to be consistently applied in an annotation protocol (Section 3).
@@
One of our primary goals is to generate QA pairs that may enhance users’ comprehension of a text going beyond its simplified version Fok et al. (2023). To this end, we pose two requirements for the QA pairs:
@@
Figure 2 illustrates the different elements and challenging nature of the task. To identify information loss, models cannot rely on lexical overlap, ordering, or other surface-level properties.
Original ($X_{\text{src}}$): These results indicate that acute/chronic endurance ($E_{1}$) arm-cranking with EMS applied to the lower limbs improves the brachial artery endothelial function ($E_{2}$) more markedly than the same exercise without EMS. Simplified ($X_{\text{tgt}}$): The study concluded that doing the arm-cranking exercise with EMS on the lower body can improve arm function ($R_{2}$) more than without it. $Q_{1}$: On what timeframes does EMS improve training? $A_{1}$: It helps in single (acute) or repeated applications (chronic). $Q_{2}$: How did the researchers measure how well EMS works? $A_{2}$: They measure to what extent the main artery of the arm widens, which is called flow-mediated vasodilation. $Q_{1}^{\prime}$ (↯): Did EMS improve training in acute and chronic applications?
Figure 2: Example with a Deletion (“acute/chronic”) and an Oversimplification (“improve arm function” is too broad given that EMS improves “artery function”). These give rise to two QA pairs ($Q_{1}$ and $Q_{2}$) which fulfill the Readability and Givenness constraints. For contrast, $Q_{1}^{\prime}$ violates (↯) givenness. $Q_{1}$ is likely more natural to lay readers because it could be asked without having seen the original text (no presupposition that the study looked at short-term and long-term effects).
@@ 3 Data Collection
(a) Distribution of QA pairs over RCT abstract sections. Some QAs are localized in multiple sections (multi.), in a rare section (other) or their abstracts do not have a section structure (19.2% of abstracts; no secs.).
(b) Cosine similarity of annotators’ questions ($Q$) by whether their evidence spans ($E$) have no/partial/complete overlap.
(c) Number of QA pairs per sentence for a random selection of 25 documents. The x-axis are sentence indices.
Figure 3: Distribution of information loss. Humans produce a similar distribution of questions by section (a), but the questions differ in their localization (c). A similar localization results in more similar questions (b). Comparing humans to models, we see differences where questions are localized, and by extension also what they are about.
@@ 3.1 Simplification Corpus
We focus on medical texts, which provide an important and illustrative case where lay individuals may be keenly interested in understanding newly published science. Specifically, we consider abstracts of medical journal articles that describe the conduct and findings of RCTs. RCTs are the foundation of evidence-based medicine that informs patient care Sackett (1998). But those articles use technical language, which is effectively inaccessible to lay patients Zuccala (2010); August et al. (2023). This means that individuals may already interact with LLMs to simplify such texts. But automated simplification of RCTs may result in information loss, which risks readers misinterpreting findings.
To better quantify information loss in LLM-simplifications, we create a parallel simplification corpus with GPT-4 (gpt-4-0613). [^3] No curated parallel corpus exists for RCT simplification. On the related task of plain language summarization of full RCTs, Shaib et al. (2023) observed low error rates (<10% of summaries affected), but we do not yet have a quantitative understanding of information loss in LLM-simplifications. We sample RCT abstracts from the Evidence Inference v2.0 dataset DeYoung et al. (2020b); Lehman et al. (2019), originally sourced from PubMed. Table A.1 provides the simplification prompt and example outputs.
@@ 3.2 Annotation Process
The process for annotating InfoLossQA is as follows. First, annotators examine both the original and simplified versions of the text. Next they identify and highlight information loss spans. For each highlighted instance, annotators compose a QA pair. We used the Thresh toolkit for annotation Heineman et al. (2023). See Appendix B for annotation guidelines and interface screenshots.
@@
Technical texts like RCTs have a long tail of information that can be lost in simplification (e.g., numerical results and significance values). Annotating all information loss is intractable, and the resulting QAs would likely overwhelm readers. Therefore, we grounded annotation in a scenario. Annotators were asked to imagine a patient who could benefit from better understanding the RCT at hand, and to focus on information that may be salient in this scenario.
@@
We hired three expert annotators who are native English speakers, majoring in linguistics and experienced in annotating medical texts. To ensure high annotation quality, we conducted a pilot phase with written feedback and held frequent review meetings to discuss questions. Annotators were compensated \$15/hr. [^4] Higher than local minimum wage at the time of study. The median duration per document was 22 minutes.
@@
We collected annotations on 104 abstract-simplification pairs (average number of tokens: $|X_{\text{src}}|=312$; $|X_{\text{tgt}}|=271$) yielding 1,000 questions and answers (average number of tokens: $|Q|=11.4$; $|A|=30.3$). [^5] Tokenizing with nltk Bird et al. (2009). Each annotator wrote on average between 4.1 and 5.7 QA pairs (i.e., information loss instances) per document. We observe that oversimplification is more prevalent (74.1% of questions) in this corpus than complete deletion of information (25.9% of questions). Additional statistics are given in Table A.2.